Find the magnitude and direction of the vector represented by each of the following pairs of components:


(a) 𝐴𝑥 = 3.0 𝑐𝑚, 𝐴𝑦 = −4.0 𝑐𝑚;


(b) 𝐴𝑥 = −5.0 𝑚, 𝐴𝑦 = −12.0 𝑚;


(c) 𝐴𝑥 = −2.0 𝑘𝑚, 𝐴𝑦 = 3.0 𝑘𝑚.



1
Expert's answer
2022-04-27T13:34:42-0400
"A=\sqrt{A_x^2+A_y^2}"

"\theta=\tan^{-1}\frac{A_y}{A_x}"

(a)

"A=\sqrt{(3.0)^2+(-4.0)^2}=5.0\:\rm cm"

"\theta=\tan^{-1}\frac{-4.0}{3.0}=307^\circ"

(b)

"A=\sqrt{(-5.0)^2+(-12.0)^2}=13\:\rm m"

"\theta=\tan^{-1}\frac{-12.0}{-5.0}=247^\circ"

(c)

"A=\sqrt{(-2.0)^2+(3.0)^2}=3.6\:\rm km"

"\theta=\tan^{-1}\frac{3.0}{-2.0}=124^\circ"